5. Consider the flux across the line segment from (1,2) to (3,2) (a) Use vectors and following the process in the previous page to compute the flux across the line segment Include umi Segment Length = Total Flux ix- b) Explain why your answer to part (a) makes sense intuitively 6. Now consider the flux across the line segment from (1,2) to (2,1) (a) Uac vortors to comp the ux ars the l Include mits.
Answer #1

here flux is how much \vec{A} pass through that line element.

flux \phi=\vec{A}.\vec{dl}=\vec{A}.\hat{n}dl where \vec{dl} is the line and \hat{n} is the unit vector showing the direction which is \perp to the line \vec{dl} .

(a) let \vec{F}=F_0\hat{i}

line element di = (3-1)i + (2-2)j = 2;

\left | \vec{dl} \right |=\sqrt{2^2}=2

lets consider any vector \vec{K}=a\hat{i}+b\hat{j} (dont need z-component as we are in x-y plane) which is \perp to \vec{dl} . then surely the unit vector \hat{n}=\frac{\vec{K}}{\left | \vec{K} \right |}

since \vec{K} is \perp to \vec{dl} , the scaler product \vec{K}.\vec{dl}=0

\vec{K}.\vec{dl}=0

\left ( a\hat{i}+b\hat{j}\right ).2\hat{i}=0

2a=0

\therefore a=0

let the other component be b=\lambda_1 where \lambda_1 are some constant.

so, \vec{K}=\lambda_1\hat{j} and 1

hence \hat{n}=\frac{\vec{K}}{\left | \vec{K} \right |}=\frac{\lambda_1\hat{j}}{\lambda_1}=\hat{j}

total flux= \vec{F}.\hat{n}\left | \vec{dl} \right |=F_0(\hat{i}.\hat{j})\times2=0

It is also very clear from the diagram that there is no \vec{F} pasing through the line segment. so only fair result is that flux should zero.

(b) \vec{dl}=(2-1)\hat{i}+(1-2)\hat{j}=\hat{i}-\hat{j}

\left | \vec{dl} \right |=\sqrt{2}

again let any vector \vec{K}=a\hat{i}+b\hat{j} which is \perp to \vec{dl} . then we have

\vec{K}.\vec{dl}=0

\left ( a\hat{i}+b\hat{j} \right ).(\hat{i}-\hat{j})=0

a-b=0

\therefore a=b

let a=b=\lambda_2 where \lambda_2 is a constant

then \vec{K}=\lambda_2(\hat{i}+\hat{j}) and \left | \vec{K} \right |=\sqrt{2}\lambda_2

unit vector \hat{n}=\frac{\vec{K}}{\left | \vec{K} \right |}=\frac{\hat{i}+\hat{j}}{\sqrt2}

then total flux across the line, F_0\hat{i}.\hat{n}dl=F_0\hat{i}\times\frac{\hat{i}+\hat{j}}{\sqrt2}\times\sqrt2=F_0